1/2b+4=1/8b+88. how do you do these one? cause is irritating me.
Answer by checkley77(12844)

1/2b+4=1/8b+88
b/2+4=b/8+88
b/2-b/8=88-4
(4b-b)/8=84
3b/8=84
3b=84*8
3b=672
b=672/3
b=224 ans.
Proof:
224/2+4=224/8+88
112+4=28+88
116=116